OpenStudy (anonymous):

Help please! (: medal + fan .Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

Is there choices?

OpenStudy (anonymous):

unfortunately not..

OpenStudy (anonymous):

5 adult tickets and 5 childrens tickets 6x5=30 10x5=50, 50+30=80

OpenStudy (anonymous):

If you multiply 6 and 5 you get 30. Which is $30 for the children. 10 x __ = 80 80 - 30 = 50 Which is 10 x 5. SO 5

OpenStudy (anonymous):

Well thanks datass.

OpenStudy (anonymous):

No problem. :p

OpenStudy (anonymous):

5 adults 10a+6(5)=80 10a+30=80 10a-30+30=80-30 10a=50 10a/10=50/10 a=5 there are six children, 30 must be spent, 50 is left, since each adult costs 10, 5 there are 5 adult ticket

OpenStudy (anonymous):

Good Job!

OpenStudy (anonymous):

I did not know if that was correct or not